# A Survey on Vision-Language-Action Models for Embodied AI

### v2

@@ III-A2 Pretrained Visual Representations
RPT [57] ViT [30, 14] TFM MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x_{i}$: token from any modality; $y,z$: other modalities [SC] Real-Mani (Franka): stack, pick, pick from bin
GR-1 [58] ViT-MAE TFM MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y\right)$ $x_{i}$: image patch; $y$: language instruction Ego4D Sim-Mani: CALVIN
[…]
Table II: Dynamics learning methods for VLAs. $f(\cdot)$ is the dynamic model. Fwd: forward. Inv: Inverse.
Model V-Encoder Objective Formula Notation Robotic Tasks
Vi-PRoM [71] ResNet [29] Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$ predicts the frame index in original seq. $x$, given a shuffled seq. $x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open and close drawer/door
@@
III-A3 Dynamics Learning
Dynamics learning encompasses objectives aimed at imbuing the model with an understanding of forward or inverse dynamics. Forward dynamics involves predicting the subsequent state resulting from a given action, whereas inverse dynamics entails determining the action required to transition from a previous state to a known subsequent state. Some research approaches also frame these objectives as reordering problems for shuffled state sequences. While forward dynamics models closely relate to world models, this subsection focuses specifically on works that leverage dynamics learning as an auxiliary task to enhance the performance of the primary robot task. We compare them in Table II.
@@ III-A3 Dynamics Learning
Perception-Action Causal Transformer (PACT) [76] introduces a pretraining objective aimed at modeling state-action transitions. PACT receives sequences of states and actions as input, and predicts each state and action token in an autoregressive manner. This pretrained model serves as a dynamics model, which can then be finetuned for various downstream tasks such as localization, mapping, and navigation.
@@
Video Pretraining (VPT) [78] proposes a method that harnesses unlabeled internet data to pretrain a foundational model for the game of Minecraft. The approach begins by training an inverse dynamics model using a limited amount of labeled data, which is then utilized to label internet videos. Subsequently, this newly auto-labeled data is employed to train the VPT foundation model through behavior cloning. This methodology follows semi-supervised imitation learning. As a result of this process, the model demonstrates human-level performance across a multitude of tasks.
III-A4 Classical World Model
@@
Dreamer [79] employs three primary modules to construct a latent dynamics model: the representation model, responsible for encoding images into latent states; the transition model, which captures transitions between latent states; and the reward model, predicting the reward associated with a given state. Under the actor-critic framework, Dreamer utilizes an action model and a value model to learn behavior through imagination by propagating analytic gradients through the learned dynamics. Building upon this foundation, DreamerV2 [80] introduces a discrete latent state space along with an improved objective. DreamerV3 [81] extends its focus to a broader range of domains with fixed hyperparameters. DayDreamer [82] applies this method to physical robots performing real-world tasks.
Masked World Model (MWM) [83] innovates by modifying the vision encoder of DreamerV2 to a hybrid composition of Convolutional Neural Network and Vision Transformer. In training this novel Convolutional-ViT vision encoder, MWM draws inspiration from the approach proposed in MAE [64]. Through the incorporation of an auxiliary reward prediction loss, the resulting learned latent dynamics model exhibits a notable performance improvement across various visual robotic tasks.
Iso-Dream [84] introduces two key enhancements to the Dreamer framework. The first enhancement focuses on optimizing inverse dynamics by decoupling controllable and noncontrollable dynamics. The second enhancement involves optimizing agent behavior based on the decoupled latent imaginations. This approach is particularly advantageous, as the noncontrollable state transition branch can be rolled out independently of actions, yielding benefits for long-horizon decision-making processes.
Transformer-based World Model (TWM) [85] shares the same motivation as Dreamer but adopts a different approach by constructing a world model based on the Transformer-XL architecture [86]. This Transformer-based world model trains a model-free agent in latent imagination by generating new trajectories. Additionally, TWM suggests a modification to the balanced KL divergence loss from DreamerV2 and introduces a novel thresholded entropy loss tailored for the advantage actor-critic framework.
@@
IRIS [87] employs a GPT-like autoregressive Transformer as the foundation of its world model, with a VQ-VAE serving as the vision encoder. Subsequently, a policy is trained using imagined trajectories, which are unrolled from a real observation by the world model, akin to TWM’s approach.
SWIM [88] advocates for the use of human videos in training a world model due to the availability of large-scale human-centric data. However, acknowledging the substantial gap between human and robot data, SWIM addresses this disparity by grounding actions in visual affordance maps. This approach involves inferring target poses based on the affordance maps, facilitating an effective transfer of knowledge from human data to enhance robot control.
@@
III-A5 LLM-induced World Model
DECKARD [90] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs, specifically tailored for the task of item crafting in Minecraft. DECKARD iterates between two phases: in the Dream phase, it samples a subgoal guided by the AWM; in the Wake phase, DECKARD executes the subgoal and updates the AWM through interactions with the game. This guided approach, centered around AWMs, enables DECKARD to achieve markedly faster item crafting compared to baselines lacking such guidance.
RAP [93] repurposes an LLM to act as both a policy that predicts actions and a world model that provides the state transition distribution. Unlike previous chain-of-thought prompting methods, RAP incorporates Monte Carlo Tree Search (MCTS) to enable principled planning, allowing the LLM to build a reasoning tree incrementally. This reasoning strategy helps RAP find a high-reward path that balances exploration and exploitation. Tree-Planner [94] improves efficiency by prompting the LLM only once to generate diverse paths within an action tree.
LLM-MCTS [95] builds upon RAP but extends the problem setting to Partially Observable Markov Decision Processes (POMDPs). It proposes using the LLM as both a world model and a heuristic policy within an MCTS planning algorithm. As a world model, the LLM generates the initial belief of the current state; as a policy, it serves as a heuristic to guide action selection. By leveraging its commonsense knowledge, the LLM reduces the search space of MCTS, thereby improving search efficiency.
E2WM [96] is an LLM training paradigm that focuses on collecting embodied experiences from world models using MCTS. Unlike prior works, E2WM treats simulators as world models, as they emulate real-world physical interactions, and MCTS is used solely for data collection. By fine-tuning LLMs with this embodied knowledge, the models gain enhanced embodied abilities, such as plan generation, activity recognition, and tracking.
@@ III-A5 LLM-induced World Model
III-A6 Reasoning
Reasoning has become an important capability in LLMs, as pioneered by chain-of-thought (CoT) methods [98, 99]. Recent OpenAI o1 models [100] further demonstrate its potential in building more powerful AI systems. A similar trend has also emerged in the field of embodied AI. EmbodiedGPT [101] uses CoT to enhance embodied planning, and ThinkBot [102] applies it to recover missing action descriptions. RAT [103] combines retrieval-augmented generation and CoT to elicit context-aware reasoning, improving long-horizon generation. ReAct [104] interleaves verbal reasoning traces and actions to address diverse decision-making tasks. Tree-Planner [94] employs a Tree-of-Thoughts approach. Text2Motion [105] uses symbolic and geometric reasoning to verify the feasibility of an LLM-generated task plan. Reflexion [106] proposes a novel verbal reinforcement learning framework, replacing weight updates in RL models with linguistic feedback.
Pros and Cons. Pretrained visual representations underscore the importance of the vision encoder, as visual observation plays a crucial role in perceiving the current state of the environment. Consequently, it sets an upper bound on the performance of the overall model. In VLAs, general vision models are pretrained using robot or human data to enhance their capabilities in tasks such as object detection, affordance map extraction, and even vision-language alignment, which are essential for robotic tasks. In contrast, dynamics learning focuses on understanding the transition between states. This involves not only mapping visual observations to good state representations but also comprehending how different actions lead to different states, and vice versa. Existing dynamics learning methods typically aim to capture the relationship between states and actions using simple masked modeling or reordering objectives. On the other hand, a world model aims to completely model the dynamics of the world, enabling the robot models to roll out states multiple steps into the future based on the current state, facilitating better prediction of optimal actions. Therefore, while a world model is more desired, it is also more challenging to achieve.
@@ III-A6 Reasoning
Table III: Low-level control policies. We also include some non-VLA models as they are closely related, marked in gray text. BC: behavior cloning (action type, cont/disc: continuous/discrete). TFM: Transformer. Xattn: cross-attention. Concat: concatenation. $p$/$s$: prompt/state vision encoder. (SC): self-collect data. LMP: Latent Motor Plans [107]. MPC: model predictive control.
@@
CLIPort [13] CLIP-ResNet50, Transporter-ResNet [108] CLIP-GPT Hadamard LingUNet [109] BC ($\textbf{SE}(2)$) (SC) Sim: Ravens [108]
MCIL [110] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP + (SC) Sim: 3D Playroom environment
@@
MCIL CNN Sentence-BERT [113] LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
@@
CLIP-GPT Concat STORM [116] Max Likelihood Est. (cost map) (SC) Sim & Real (Franka): pick, place
Interactive Language [117] ResNet [29] CLIP-GPT Xattn TFM BC (cont) Language-Table (SC) Sim & Real (xArm6): rearrangement
Hiveformer [118] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data [119] Sim: RLBench
PerAct [120] 3D CNN CLIP-GPT Concat PerceiverIO [121] 3D affordance RLBench data [119] Sim: RLBench; Real (Franka): put marker, place food, etc.
Act3D CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
RVT, RVT-2 [122, 123] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data [119], (SC) Sim: RLBench; Real (Franka): stack, press, place
RoboPoint [124] ViT-L/14 Vicuna-V1.5 13B [125] Concat - 2D affordance (project to 3D) (SC) Real (Franka): pick, place
Gato [50] ViT Sent.Piece [126] Concat TFM BC (cont & disc) (SC) Sim & Real (Sawyer): RGB-stacking
@@
VIMA [129] ViT, Mask R-CNN [74] T5 [10] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data (SC) Sim (Ravens): VIMA-Bench
BC-Z [130] ResNet18 ($p,s$) [29] USE FiLM [131] MLP BC (cont) (SC) Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
RT-1 [3] EfficientNet [132] USE FiLM TFM BC (disc) Kitchen (SC) Real (Everyday Robots): pick-place, move, knock
MOO [129] OWL-ViT ($p$) [133], EfficientNet ($s$) USE FiLM TFM BC (disc) (SC) Real (Everyday Robots): pick, move near, knock, place upright, place into
Q-Transformer [134] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1 data Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
RT-Trajectory [135] EfficientNet - - TFM BC (disc) (SC) Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
ACT [136] ResNet18 - - CVAE-TFM BC (cont, action chunking) (SC) with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
MT-ACT [137] CNN N/A FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
RT-2 [2] ViT-22B [138],
ViT-4B [139]
PaLI-X [140],
PaLM-E
Concat Symbol-tuning [141] BC (disc), Co-fine-tuning VQA + Kitchen Real: RT-1 evaluation tasks, Language-Table
RT-H [142] ViT-22B [138] PaLI-X 55B [140] Concat Symbol-tuning BC (disc) Diverse + Kitchen Real: Kitchen, Diverse
RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE (SC) Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
OpenVLA [144] DINOv2 [145],
SigLIP [146]
LLaMA 2 [147] (Prismatic-7B VLM [148] Concat Symbol-tuning BC (disc) OXE, DROID [149] Real: BridgeV2, RT-1 evaluation tasks
RoboFlamingo [46] CLIP-ViT-L/14 LLaMA [150], GPT-NeoX [151], MPT [152] Xattn LSTM, TFM, MLP BC (cont) CALVIN data Sim: CALVIN
@@
GPT-4 [157] Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
UniPi [158] Imagen video [159] T5-XXL [10] Xattn CNN Inverse dynamics PF (SC), Ravens, BridgeV1 Sim: Painting Factory (PF) [160], Ravens
Diffusion Policy [161] ResNet18 [29] - - UNet,
@@
DDPM [162] (SC) Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
DP3 [163] DP3 Encoder (Model design follows Diffusion Policy) DDIM (SC) Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
SUDD [164] ResNet18 [29] CLIP-GPT Concat UNet,
@@
Octo [165] CNN T5-base [10] Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
3D Diffuser Actor [166] (Model design follows Act3D) DDPM RLBench, CALVIN, (SC) Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
MDT [167] CLIP-ViT-B/16 ($p$), Voltron / ResNet18 ($s$) CLIP-GPT Concat DiT [168] DDIM [169] CALVIN, LIBERO, Kitchen (SC) Sim: CALVIN, LIBERO [170]; Real (Franka): pick, push, open, close, etc.
@@
III-B Low-level Control Policy
Through the integration of an action decoder with perception modules, such as vision encoders and language encoders, a control policy network is formed to execute language instructions in both simulated and real-world environments. The diversity among control policy networks arises from the selection of individual modules and the integration strategy. This section explores various approaches to designing a low-level control policy. Based on the backbone of their action decoder, control policies can be broadly categorized into three types: non-Transformer (§III-B1), Transformer-based (§III-B2), and LLM-based (§III-B8). Other approaches focus on specific aspects of control policies, such as multi-modal instructions (§III-B3), 3D Vision (§III-B4), diffusion-based action generation (§III-B5), and motion planning (§III-B7). In Table III, we compare the architectural details of these control policies.
III-B1 Early Non-Transformer Control Policies

### v4

@@ III-A2 Pretrained Visual Representations
RPT [42] ViT [17, 41] MAE $\sum_{i=1}^{N}-\log P\left(x_{i}\mid x_{\neq i},y,z,\dots\right)$ $x,y,z$: three distinct modalities Real-Mani (Franka): stack, pick, pick from bin
DINOv2 [49] ResNet, ViT Self-distillation $\sum_{x}\sum_{x^{\prime}\neq x}H(P_{t}(x),P_{x}(x^{\prime}))$ $x,x^{\prime}$: image views; $H()$: cross-entropy; $P_{t},P_{s}$: teacher, student (Used by OpenVLA [50], ReKep [51])
I-JEPA [52] ViT JEPA $\frac{1}{M}\sum_{i=1}^{M}\sum_{j}\left\lVert x_{j}(i)-y_{j}(i)\right\rVert_{2}^{2}$ $x_{j}(i)$: $j$-th masked patch embedding of block $i$; $y_{j}(i)$: unmasked patch embedding -
Theia [53] ViT-T/S/B Distillation (Distillation of vision foundation models, including ViT, CLIP, SAM, DINOv2, and Depth-Anything.) Sim-Mani&Navi: CortexBench [47]; Real-Mani: pick, place, open door/drawer
Voltron [48] introduces a pretraining objective by incorporating language conditioning and language generation into the MAE objective. Employing an encoder-decoder Transformer structure, the pretraining alternates between language-conditioned masked image reconstruction and language generation from masked images. This enhances the alignment between language and vision modalities.
VC-1 [47] conducts an in-depth examination of prior PVRs and introduces an enhanced PVR model by systematically exploring optimal ViT configurations across diverse datasets. Additionally, they perform a comprehensive comparative analysis of their model against previous methods on various manipulation and navigation datasets, shedding light on the critical factors that contribute to the improvement of PVR. Another study [54] also compares previous PVRs obtained under supervised learning or self-supervised learning.
DINOv2 [49] proposes a new self-supervised training paradigm for PVRs that achieves performance beyond that of MAE. It employs a self-distillation framework in which the teacher and student networks receive different views of the same image and match their encoded representations. The student network is updated using SGD, while the teacher network is maintained as the EMA of the student network.
I-JEPA [52] is motivated by the joint-embedding predictive architectures proposed by [55]. It constructs a “primitive” internal world model by comparing the embeddings of patches. Unlike DINO, which uses cropped images, I-JEPA employs masked patches. Moreover, it differs from MAE because it is a non-generative approach.
Theia [53] proposes distilling various vision foundation models into a single model. By fusing their information from segmentation, depth, semantics, etc., it outperforms previous PVRs while requiring less data and a smaller model size.
III-A3 Video Representations
Videos are simple sequences of images and can be represented by concatenating the usual PVRs of each frame. However, their multi-view nature enables a variety of unique representation techniques beyond those mentioned above, such as time contrastive learning and MAE. NeRF can be extracted from videos and contains rich 3D information for robot learning, as exemplified by F3RM [56] and 3D-LLM [57]. The recent 3D Gaussian Splatting (3D-GS) method has been gaining popularity in computer vision as well as in embodied AI. 3D Gaussians can be deformed through physics-based simulation for generative dynamics, as proposed in PhysGaussian [58], or they can serve as 3D representations for VLMs, as enabled by UniGS [59]. Additionally, many videos contain audio, which can provide important cues for robot policies [60].
III-A4 Dynamics Learning
Dynamics learning encompasses objectives aimed at endowing the model $f(\cdot)$ with an understanding of forward or inverse dynamics. Forward dynamics involves predicting the subsequent state resulting from a given action, whereas inverse dynamics entails determining the action required to transition from a previous state to a known subsequent state:
$\begin{split}\text{Forward dynamics: }&\hat{s}_{t+1}\leftarrow f_{\text{fwd}}(s_{t},a_{t}),\\ \text{Inverse dynamics: }&\hat{a}_{t}\leftarrow f_{\text{inv}}(s_{t},s_{t+1}).\end{split}$ (1)
Some approaches also frame these objectives as reordering problems for shuffled state sequences. Some forward dynamics methods closely resemble the image or video prediction pretraining used in PVRs. We compare them in Table II.
Table II: Dynamics learning methods for VLAs. $f(\cdot)$ is the dynamic model. Fwd, inv: forward & inverse dynamics.
Model V-Encoder Type Objective Notation Robotic Tasks
Vi-PRoM [61] ResNet Temporal dynamics $\sum_{i=1}^{T}\operatorname{CE}(i,f(x_{i}\mid x^{\prime}))$ $f(\cdot)$: predicts frame index in raw seq. $x$ given shuffled seq. $x^{\prime}$ Sim-Mani: Meta-World, Franka Kitchen; Real-Mani: open and close drawer/door
@@ III-A4 Dynamics Learning
GR-1 [69] ViT-MAE Forward dynamics $\sum_{i=1}^{T}\operatorname{MSE}\left(s_{t+1},f_{\text{fwd}}(s_{t},a_{t})\right)$ $s_{t}$, $a_{t}$: state, action Sim-Mani: CALVIN
@@
PACT [66] introduces a pretraining objective aimed at modeling state-action transitions. It receives sequences of states and actions as input, and predicts each state and action token autoregressively. This pretrained model serves as a dynamics model, which can then be finetuned for various downstream tasks such as localization, mapping, and navigation.
@@
VPT [68] proposes a video pretraining method that harnesses unlabeled internet data to pretrain a foundational model for the game of Minecraft. The approach begins by training an inverse dynamics model using a limited amount of labeled data, which is then utilized to label internet videos. Subsequently, this newly auto-labeled data is employed to train the VPT foundation model through behavior cloning. This methodology follows semi-supervised imitation learning. As a result of this process, the model demonstrates human-level performance across a multitude of tasks.
GR-1 [69] introduces video prediction pretraining tailored for a GPT-style model. The ability to anticipate future frames aligns with forward dynamics learning, contributing to more accurate action prediction.
III-A5 World Models
A world model $P(\cdot)$ encodes commonsense knowledge about the world and predicts the future state for a given action [55]:
$\begin{split}\hat{s}_{t+1}\sim P(\hat{s}_{t+1}|s_{t},a_{t}).\end{split}$ (2)
It enables model-based control and planning for embodied agents, as they can search for an optimal action sequence in imaginary space before executing any real actions. Although forward dynamics learning also attempts to predict the next state, it is usually treated as a pretraining task or an auxiliary loss to enhance the RL-Transformer-based action decoder for the primary robotics tasks, rather than serving as a standalone module. Additionally, new embodied experiences can be sampled from visual world models that explicitly generate images or videos of future states.
@@ III-A5 World Models
Dreamer [70] employs three primary modules to construct a latent dynamics model: a representation model, responsible for encoding images into latent states; a transition model, which captures transitions between latent states; and a reward model, predicting the reward associated with a given state. Under the actor-critic framework, Dreamer utilizes an action model and a value model to learn behavior through imagination by propagating analytic gradients through the learned dynamics. Building upon this foundation, DreamerV2 [71] introduces a discrete latent state space along with an improved objective. DreamerV3 [72] extends its focus to a broader range of domains with fixed hyperparameters. DayDreamer [73] applies this method to physical robots performing real-world tasks.
@@
IRIS [74] employs a GPT-like autoregressive Transformer as the foundation of its world model, with a VQ-VAE serving as the vision encoder. Subsequently, a policy is trained using imagined trajectories, which are unrolled from a real observation by the world model. TWM [75] also investigates the application of Transformers in building world models.
III-A6 LLM-induced World Models
LLMs encompass a wealth of commonsense knowledge about the world, prompting many approaches to leverage that knowledge for improving VLAs.
DECKARD [76] prompts LLM to generate abstract world models (AWMs) represented as directed acyclic graphs, specifically tailored for the task of item crafting in Minecraft. DECKARD iterates between two phases: in the Dream phase, it samples a subgoal guided by the AWM; in the Wake phase, DECKARD executes the subgoal and updates the AWM through interactions with the game. This guided approach enables DECKARD to achieve markedly faster item crafting compared to baselines lacking such guidance.
RAP [79] repurposes an LLM to act as both a policy that predicts actions and a world model that provides the state transition distribution. Unlike previous chain-of-thought prompting methods, RAP incorporates Monte Carlo Tree Search (MCTS) to enable structured planning, allowing the LLM to build a reasoning tree incrementally. This reasoning strategy helps RAP find a high-reward path that balances exploration and exploitation. Tree-Planner [80] improves efficiency by prompting the LLM only once to generate diverse paths within an action tree.
LLM-MCTS [81] builds upon RAP but extends the problem setting to POMDPs. As a world model, the LLM generates the initial belief of the current state; as a policy, it serves as a heuristic to guide action selection. By leveraging its commonsense knowledge, the LLM reduces the search space of MCTS, thereby improving search efficiency.
III-A7 Visual World Models
Unlike LLM-induced world models, which are in textual form, visual world models generate images, videos, or 3D scenes of future states—aligning more closely with the physical world. They can further be utilized to generate new trajectories. This direction has been gaining increasing attention since Sora demonstrated world-simulation capabilities [82], as investigated by a dedicated survey [83].
@@ III-A7 Visual World Models
UniSim [86] builds a generative model based on real-world interaction videos. It is capable of simulating visual outcomes for both high-level and low-level actions, which can then be leveraged as new experiences for training embodied agents. E2WM [87] even treats existing simulators as world models to collect embodied experiences via MCTS.
III-A8 Reasoning
Reasoning has become a key capability of LLMs, as demonstrated by chain-of-thought (CoT) methods [88, 89]. In embodied AI, researchers are exploring how to leverage CoT reasoning to refine the decision-making process.
ThinkBot [90] applies it to recover missing action descriptions in sparse human instructions. RAT [91] combines retrieval-augmented generation and CoT to elicit context-aware reasoning, improving long-horizon generation. ReAct [92] interleaves verbal reasoning traces and actions to allow for diverse types of reasoning. Tree-Planner [80] employs a tree-of-thoughts approach. ECoT [93] re-trains OpenVLA [50] using reasoning chain datasets, enabling embodied CoT reasoning that further enhances its performance.
Pros and Cons.
PVRs
Although time contrastive learning with videos and text-guided pretraining methods, such as CLIP, provide image-level information, they lack the pixel-level details offered by MAE-based self-supervised learning methods. Pixel-level information contains rich details—including segmentation masks, object positions, and depth estimation—that are generally more useful for robot manipulation tasks requiring high precision, as demonstrated by VC-1’s comparison [47]. DINOv2 learns both pixel- and image-level features by combining masked image modeling with a momentum encoder and multi-crop augmentation, with benefits on downstream robotic tasks evidenced by OpenVLA [50]. I-JEPA focuses on patches in the representation space and, as a result, captures low-level image features more effectively than view-invariance methods such as DINO. Theia distills various off-the-shelf vision foundation models into a single model that surpasses isolated individual models, as demonstrated by comprehensive evaluations against most existing PVRs in robot learning [53].
Forward & inverse dynamics
Forward dynamics learning is generally more challenging than inverse dynamics learning because predicting future states is more complex than predicting past actions. Consequently, the difficulty of forward dynamics often leads to greater performance improvements, as demonstrated in SMART. However, inverse dynamics models can be used to generate action labels for datasets that contain only states, such as raw robot manipulation videos.
World Models & Reasoning
Although both world models and reasoning methods can be applied to low-level control policies and high-level task planners, current approaches remain distinct. World models are predominantly used to interact with control policies because they excel at generating the immediate next state given low-level actions. In contrast, CoT-based reasoning methods focus on task planning since they express thought chains in text, making them well-suited for refining text-based task plans.
Figure 4: Illustration of a hierarchical robot policy. The high-level task planner decomposes the user instruction into subtasks, which are then executed step by step by the low-level control policy.
@@ III-A8 Reasoning
Table III: Low-level control policies. We also include some non-VLA models as they are closely related, marked in gray text. BC: behavior cloning (action type, cont/disc: continuous/discrete). TFM: Transformer. Xattn: cross-attention. Concat: concatenation. $p$/$s$: prompt/state vision encoder. [SC]: self-collect data. MPC: model predictive control.
@@
CLIPort [43] CLIP-ResNet50, Transporter-ResNet CLIP-GPT Hadamard LingUNet BC ($\textbf{SE}(2)$) [SC] Sim: Ravens [94]
MCIL [95] Custom-CNN USE LMP RNN LMP RNN MCIL Play-LMP, [SC] Sim: 3D Playroom environment
@@
MCIL CNN Sentence-BERT LMP TFM LMP RNN MCIL CALVIN data Sim: CALVIN
@@
CLIP-GPT Concat STORM Max Likelihood Est. (cost map) [SC] Sim & Real (Franka): pick, place
Interactive Language [100] ResNet CLIP-GPT Xattn TFM BC (cont) Language-Table [SC] Sim & Real (xArm6): rearrangement
Hiveformer [101] UNet CLIP-GPT Concat TFM, UNet BC (cont) RLBench data Sim: RLBench [102]
PerAct [103] 3D CNN CLIP-GPT Concat PerceiverIO 3D affordance RLBench data Sim: RLBench; Real (Franka): put marker, place food, etc.
Act3D [104] CLIP-ResNet50 CLIP-GPT Xattn TFM 3D affordance (voxel) RLBench, (BC) Sim: RLBench; Real (Franka): reach target, wipe, stack, etc.
RVT, RVT-2 [105, 106] CLIP-ResNet50 CLIP-GPT Concat TFM 2D affordance (project to 3D) RLBench data, [SC] Sim: RLBench; Real (Franka): stack, press, place
RoboPoint [107] ViT-L/14 Vicuna-V1.5 13B [108] Concat - 2D affordance (project to 3D) [SC] Real (Franka): pick, place
Gato [36] ViT Sent.Piece Concat TFM BC (cont & disc) [SC] Sim & Real (Sawyer): RGB-stacking
@@
VIMA [111] ViT, Mask R-CNN [64] T5 [112] Xattn TFM BC ($\textbf{SE}(2)$) VIMA-Data [SC] Sim (Ravens): VIMA-Bench
BC-Z [113] ResNet18 ($p,s$) USE FiLM MLP BC (cont) [SC] Real (Everyday Robots): pick-place, pick-wipe, pick-drag, grasp, push
RT-1 [114] EfficientNet [115] USE FiLM TFM BC (disc) RT-1-Kitchen [SC] Real (Everyday Robots): pick-place, move, knock
MOO [111] OWL-ViT ($p$) [116], EfficientNet ($s$) USE FiLM TFM BC (disc) [SC] Real (Everyday Robots): pick, move near, knock, place upright, place into
Q-Transformer [117] EfficientNet USE FiLM TFM TD error Auto-collect + RT-1-Kitchen Sim: pick; Real (Everyday Robots): pick, place, open/close drawer, move near
RT-Trajectory [118] EfficientNet - - TFM BC (disc) [SC] Real (Everyday Robots): pick, place, fold towel, swivel chair, etc.
ACT [119] ResNet18 - - CVAE-TFM BC (cont, action chunking) [SC] with ALOHA Sim: transfer cube, bimanual insertion; Real (ViperX, WidowX): slot battery, open cup, etc
MT-ACT [120] (RoboAgent) CNN - FiLM CVAE-TFM BC (cont, action chunking) RoboSet Real (Franka): pick, place, open/close drawers, pour, push, drag, etc.
RoboFlamingo [121] CLIP-ViT-L/14 LLaMA [122], MPT [123] Xattn LSTM, TFM BC (cont) CALVIN data Sim: CALVIN
@@
GPT-4 Code MPC - - Sim: Sapien; Real (Franka): move & avoid, set up table, close drawer, open bottle, sweep Trash
UniPi [128] Imagen video [129] T5-XXL Xattn CNN Inverse dynamics PF [SC], Ravens, BridgeV1 Sim: Painting Factory (PF), Ravens
Diffusion Policy [130] ResNet18 - - UNet,
@@
DDPM [131] [SC] Sim: Robomimic, Franka Kitchen, etc; Real (UR5, Franka): push-T, flip mug, pour sauce
DP3 [132] DP3 Encoder (Model design follows Diffusion Policy) DDIM [SC] Sim: (72 tasks); Real (Franka, Allegro hand): roll-up, dumpling, drill, pour
SUDD [133] ResNet18 CLIP-GPT Concat UNet,
@@
Octo [134] CNN T5-base Concat TFM DDPM OXE Real: BridgeV2, Stanford Coffee, Berkeley Peg Insert, etc.
3D Diffuser Actor [135] (Model design follows Act3D) DDPM RLBench, CALVIN, [SC] Sim: RLBench, CALVIN; Real (Franka): put, open, close, stack, etc.
MDT [136] CLIP-ViT-B/16 ($p$), Voltron/ResNet18 ($s$) CLIP-GPT Concat DiT [137] DDIM [138] CALVIN, LIBERO, [SC] Sim: CALVIN, LIBERO; Real (Franka): pick, push, open, close, etc.
@@
RT-2 [2] ViT-22B [140],
ViT-4B [141]
PaLI-X, PaLM-E Concat Symbol-tuning BC (disc), Co-fine-tuning VQA, RT-1-Kitchen Real: RT-1 evaluation tasks, Language-Table
RT-H [142] ViT-22B [140] PaLI-X 55B Concat Symbol-tuning BC (disc) Diverse+Kitchen Real: Diverse+Kitchen eval tasks
RT-X [143] (Models from RT-1 and RT-2) BC (disc) OXE [SC] Real: Kitchen Mani., NYU Door Open, BridgeV2, etc.
OpenVLA [50] DINOv2, SigLIP [144] Prismatic-7B [145] Concat Symbol-tuning BC (disc) OXE, DROID [146] Real: BridgeV2, RT-1 evaluation tasks
$\pi_{0}$ [147] SigLIP PaliGemma Concat Action expert Flow matching OXE, $\pi$-cross-embod. [SC] Real (general robot control): sweep, open, pack, etc.
III-B Low-level Control Policies
Through the integration of an action decoder with perception modules, such as vision encoders and language encoders, a VLA model $\pi_{\theta}$ with parameters $\theta$ is formed as a control policy to execute language instructions $p$:
$\begin{split}\hat{a}_{t}\sim\pi_{\theta}(\hat{a}_{t}|p,s_{\leq t},a_{<t}).\end{split}$ (3)
The diversity among control policy networks arises from individual modules and overall architectures. The most typical architecture is shown in Figure 1. This section explores various approaches to designing low-level control policies. Table III summarizes their technical details.
III-B1 Non-Transformer Control Policies
Prior to the adoption of Transformer models, early control policies for language-conditioned robotic tasks varied significantly in architecture.